# Deep Neural Networks: Multi-Classification and Universal Approximation

### v1

@@ 4.2.1. The induction strategy
As we already observed, we need $L_{k}=2\alpha_{k}$ layers to compress all the elements of $\mathcal{C}_{k}$. Consequently, we have shown that for any arbitrary $k\in\llbracket 0,M-1\rrbracket$ there exist $z_{k}\in\mathbb{R}^{2}$, a depth $L_{k}\geq 1$, and parameters $\mathcal{W}^{L_{k}}$ and $\mathcal{B}^{L_{k}}$ such that (4.5) and (4.6) hold. Furthermore, we can explicitly construct the parameters by following (4.11) and (4.2.2).
@@
$\phi^{L_{1}}(\mathcal{W}^{L_{1}},\mathcal{B}^{L_{1}},\mathcal{C}_{k})=z_{k},\qquad\text{for all }k\in\llbracket 0,M-1\rrbracket.$ (4.13)
@@ 4.3. Data sorting
$\phi^{L_{2}}(\mathcal{W}^{L_{2}},\mathcal{B}^{L_{2}},z_{k})=\xi_{k},\quad\text{for all }k\in\llbracket 0,M-1\rrbracket.$ (4.14)
Let $\{\beta_{\eta}\}_{\eta=0}^{M-1}$ be a sequence of positive numbers. Note that to prove (4.14), it is sufficient to show that
$\begin{split}\phi(\mathcal{W}^{\beta_{\eta}}&,\mathcal{B}^{\beta_{\eta}},z_{k})=\xi_{k},\qquad\text{for all }k\in\llbracket 0,\eta\rrbracket,\\ &\phi(\mathcal{W}^{\beta_{\eta}},\mathcal{B}^{\beta_{\eta}},z_{\eta+1})=\xi_{M-1},\end{split}$ (4.15)
@@
Projection: Let us consider $w_{3}\in\mathbb{R}^{2}$ and $b_{3}\in\mathbb{R}$ such that
@@
  • As observed in Figure 16 the hyperplane $H_{5}^{1}$ considered satisfies the conditions (4.16) and (4.17). They are satisfied whenever $\theta\in(0,\theta^{*})$, where $\theta$ is the angle between the $x-$axis and the hyperplane $H_{5}^{1}$, and $\theta^{*}$ is the angle between the $x-$axis and hyperplane containing $z_{0}^{4}$ and the farthest point of $\{z_{k}^{4}\}_{k=1}^{M-1}$ on the $x-$axis.
  • Note that, in the initial step of the proof, we iterate five times to obtain (4.15) with $\eta=0$. But given the configuration $\{z_{k}^{5}\}_{k}$, only two extra steps (specifically, the last data structuring and projection steps) are necessary to satisfy (4.15) with $\eta=1$. Therefore, to sort the first $\eta\in\llbracket 1,M-3\rrbracket$ classes, one requires $5+2\eta$ iterations.
  • By construction, the point to be sorted is always the one that is placed the furthest from the origin. Therefore, it is enough to sort the first $M-1$ points, and the point $M$ will automatically be sorted.
@@
(2) Inductive Step: Let $\eta\in\llbracket 1,M-2\rrbracket$ and $\eta_{1}:=5+2(\eta+1)$. Assume that we have sorted the first $\eta$ points of $\{z_{k}^{\eta_{1}}\}_{k=0}^{M-1}$, that is $j_{k}=k$ for $k\in\llbracket 0,\eta\rrbracket$. We will show that we can sort one extra element.
@@
Data structuring: Let $k_{1},\,k_{2}\in\llbracket\eta,M-1\rrbracket$ be such that $z_{k_{1},j_{\eta+2}-1}^{\eta_{1}}$ and $z_{k_{2},j_{\eta+2}+1}^{\eta_{1}}$ are the left and right neighborhood points of $z_{\eta+2,j_{\eta+2}}^{\eta_{1}}$, respectively (if $z_{k_{2},j_{\eta+2}+1}^{\eta_{1}}$ does not exist, it is enough to take $z_{k_{2},j_{\eta+2}+1}^{\eta_{1}}:=z_{k_{1},j_{\eta+2}}^{\eta_{1}}+1$). Thus, consider the parameters
$w_{\eta_{1}+1}^{1}$ $=1,\quad w_{\eta_{1}+1}^{2}=-1,\quad b_{\eta_{1}+1}^{1}=-\left(\frac{z_{k_{2},j_{\eta+2}+1}^{\eta_{1}}+z_{\eta+2,j_{\eta+2}}^{\eta_{1}}}{2}\right),$
$\qquad\text{and}\quad b_{\eta_{1}+1}^{2}=\frac{z_{\eta+2,j_{\eta+2}}^{\eta_{1}}+z_{k_{1},j_{\eta+2}-1}^{\eta_{1}}}{2}.$
Define $W_{\eta_{1}+1}=(w_{\eta_{1}+1}^{1},w_{\eta_{1}+1}^{2})^{\top}$ and $b_{\eta_{1}+1}=(b_{\eta_{1}+1}^{1},b_{\eta_{1}+1}^{2})^{\top}$ and set
$z^{\eta_{1}+1}_{k}=\bm{\sigma}(W_{\eta_{1}+1}z^{\eta_{1}}_{k}+b_{\eta_{1}+1}).$
We obtain that $z^{\eta_{1}+1}_{\eta+2}=(0,0)$ and $z^{\eta_{1}+1}_{\eta+1}=(a,0)$, for some $a>0$, is the farthest point in the $x-$axis. Moreover, for a decreasing sequence of positive numbers $\{a_{k}\}_{k=0}^{M-1}$, we deduce
$z^{\eta_{1}+1}_{k}=(0,a_{k}),\quad\text{for all }k\in\llbracket 0,k_{1}\rrbracket.$ (4.18)
Projection: Let $w_{\eta_{1}+2}\in\mathbb{R}^{2}$ and $b_{\eta_{1}+2}\in\mathbb{R}$ be such that
$0\leq w_{\eta_{1}+2}\cdot z_{k}^{\eta_{1}+1}+b_{\eta_{1}+2}<w_{\eta_{1}+2}\cdot z_{k+1}^{\eta_{1}+1}+b_{\eta_{1}+2}\quad\text{for all }k\in\llbracket 0,\eta\rrbracket,$ (4.19)
$w_{\eta_{1}+2}\cdot z_{k}^{\eta_{1}+1}+b_{\eta_{1}+2}<w_{\eta_{1}+2}\cdot z_{\eta+2}^{\eta_{1}+1}+b_{\eta_{1}+2}\quad\text{for all }k\in\llbracket 0,M-1\rrbracket.$ (4.20)
The assumptions about $W_{\eta_{1}+2}$ mentioned above are not restrictive. Condition (4.19) is feasible, as shown by (4.18), where the sequence $z_{k}^{\eta_{1}+1}$ is arranged on the $y-$axis for $k\in\llbracket 0,M-2\rrbracket$ as a decreasing sequence. Furthermore, (4.20) is achievable by selecting the appropriate slope of the hyperplane defined by the parameters (see Remark 4.1). We then set $z_{k,j}^{\eta_{1}+2}=\sigma(w_{\eta_{1}+2}\cdot z_{k}^{\eta_{1}+1}+b_{\eta_{1}+2}),$ and, by construction, we have that $j_{k}=k$ for $k\in\llbracket 1,\eta+1\rrbracket$ and $j_{\eta+2}=M-1$, concluding the induction.
@@ 4.4. Mapping to the respective labels
$\phi^{L_{2}}(\mathcal{W}^{L_{2}},\mathcal{B}^{L_{2}},z_{k})=\xi_{k},\quad\text{for all }k\in\llbracket 0,M-1\rrbracket.$ (4.21)
@@
$\phi^{L_{3}}(\mathcal{W}^{L_{3}},\mathcal{B}^{L_{3}},\xi_{k})=k,\quad\text{for all }k\in\llbracket 0,M-1\rrbracket.$ (4.22)
@@
(2) Inductive Step: Consider $\eta\in\llbracket 0,M-1\rrbracket$ and define $\eta_{1}:=1+2(\eta-2)$. Let us assume that $\xi_{k}^{\eta_{1}}=k$ for all $k\in\llbracket 0,\eta\rrbracket$. We will show that there exist parameters such that $\xi_{k}^{\eta_{1}+2}=k$ for all $k\in\llbracket 0,\eta+1\rrbracket$. We will proceed in two steps: data structuring and data projection.
@@
$w_{\eta_{1}+1}^{1}=\frac{2}{\xi_{\eta+1}^{\eta_{1}}-\xi_{\eta}^{\eta_{1}}},\quad w_{\eta_{1}+1}^{2}=1,\quad b_{\eta_{1}+1}^{1}=-\left(\frac{\xi_{\eta+1}^{\eta_{1}}+\xi_{\eta}^{\eta_{1}}}{\xi_{\eta+1}^{\eta_{1}}-\xi_{\eta}^{\eta_{1}}}\right),\quad\text{and}\quad b_{\eta_{1}+1}^{2}=0.$
Define $W_{\eta_{1}+1}=(w_{\eta_{1}+1}^{1},w_{\eta_{1}+1}^{2})^{\top}$ and $b_{\eta_{1}+1}=(b_{\eta_{1}+1}^{1},b_{\eta_{1}+1}^{2})^{\top}$, and consider
$\xi_{k}^{\eta_{1}+1}=\bm{\sigma}(W_{\eta_{1}+1}\xi_{k}^{\eta_{1}}+b_{\eta_{1}+1})$
so that $\xi_{\eta+1}^{\eta_{1}+1}=(1,\xi_{\eta+1}^{\eta_{1}})$ and
$\xi_{k}^{\eta_{1}+1}=(0,k),\quad\text{for all }k\in\llbracket 0,\eta\rrbracket,$
$\xi_{k}^{\eta_{1}+1}=(a_{k},c_{k}),\quad\text{for all }k\in\llbracket\eta+2,M-1\rrbracket,$
where $\{a_{k}\}_{k=\eta+2}^{M-1}$ and $\{b_{k}\}_{k=\eta+2}^{M-1}$ are two sequences of strictly increasing numbers such that $a_{k}>1$ and $c_{k}>\xi_{\eta+1}^{\eta_{1}}$ for all $k\geq\eta+2$.
Data projection. Finally, define $\xi_{k}^{\eta_{1}+2}=\sigma(w_{\eta_{1}+2}\cdot\xi_{k}^{\eta_{1}+1}+b_{\eta_{1}+2})$ with
$w_{\eta_{1}+2}=((\eta+1)-\xi_{\eta+1}^{\eta_{1}},1),\quad\text{and }\quad b_{\eta_{1}+2}=0$
so that $\xi_{k}^{\eta_{1}+2}=k$ for all $k\in\llbracket 0,\eta+1\rrbracket$.
@@
Summarising, the input-output map $\phi^{L}$ of Theorem 1.1, for $N$ points with $M$ classes, is given by the composition of the mappings $\phi_{i}^{L_{i}}$ given by (4.2), (4.13), (4.14) and (4.22) respectively, i.e.,
$\phi^{L}=(\phi_{3}^{L_{3}}\circ\phi_{2}^{L_{2}}\circ\phi_{1}^{L_{1}}\circ\phi_{0}^{L_{0}}),$

### v2

@@ 4.2.1. The induction strategy
As we already observed, we need $\hat{L}_{k}=2\alpha_{k}$ layers to compress all the elements of $\mathcal{C}_{k}$. Consequently, we have shown that for any arbitrary $k\in\llbracket 0,M-1\rrbracket$ there exist $z_{k}\in\mathbb{R}^{2}$, a depth $\hat{L}_{k}\geq 1$, and parameters $\mathcal{W}^{\hat{L}_{k}}$ and $\mathcal{B}^{\hat{L}_{k}}$ such that (4.5) and (4.6) hold. Furthermore, we can explicitly construct the parameters by following (4.11) and (4.2.2).
@@
$\phi(\mathcal{W}^{L_{1}},\mathcal{B}^{L_{1}},\mathcal{C}_{k})=z_{k},\qquad\text{for all }k\in\llbracket 0,M-1\rrbracket.$ (4.13)
@@ 4.3. Data sorting
$\phi(\mathcal{W}^{L_{2}},\mathcal{B}^{L_{2}},z_{k})=\xi_{k},\quad\text{for all }k\in\llbracket 0,M-1\rrbracket.$ (4.14)
Let $\{\beta_{\eta}\}_{\eta=0}^{M-1}$ be a sequence of positive numbers and $\{\hat{\xi}_{k}\}_{k=0}^{M-1}\subset\mathbb{R}$ a strictly increasing sequence. Note that to prove (4.14), it is sufficient to show that
$\begin{split}\phi(\mathcal{W}^{\beta_{\eta}}&,\mathcal{B}^{\beta_{\eta}},z_{k})=\hat{\xi}_{k},\qquad\text{for all }k\in\llbracket 0,\eta\rrbracket,\\ &\phi(\mathcal{W}^{\beta_{\eta}},\mathcal{B}^{\beta_{\eta}},z_{\eta+1})=\hat{\xi}_{M-1},\end{split}$ (4.15)
@@
Projection: Due Lemma 2.1 there exist $w_{3}\in\mathbb{R}^{2}$ and $b_{3}\in\mathbb{R}$ such that
@@
$\bullet$ As observed in Figure 18 the hyperplane $H_{5}^{1}$ considered satisfies the conditions (4.16) and (4.17). They are satisfied whenever $\theta\in(0,\theta^{*})$, where $\theta$ is the angle between the $x-$axis and the hyperplane $H_{5}^{1}$, and $\theta^{*}$ is the angle between the $x-$axis and hyperplane containing $z_{0}^{4}$ and the farthest point of $\{z_{k}^{4}\}_{k=1}^{M-1}$ on the $x-$axis.
$\bullet$ Note that, in the initial step of the proof, we iterate five times to obtain (4.15) with $\eta=0$. But given the configuration $\{z_{k}^{5}\}_{k}$, only two extra steps (specifically, the last data structuring and projection steps) are necessary to satisfy (4.15) with $\eta=1$. Therefore, to sort the first $\eta\in\llbracket 1,M-3\rrbracket$ classes, one requires $5+2\eta$ iterations.
$\bullet$ By construction, the point to be sorted is always the one that is placed the furthest from the origin. Therefore, it is enough to sort the first $M-1$ points, and the point $M$ will automatically be sorted.
@@
(2) Inductive Step: Let $\eta\in\llbracket 1,M-2\rrbracket$ and $\ell:=5+2(\eta+1)$. Assume that we have sorted the first $\eta$ points of $\{z_{k}^{\ell}\}_{k=0}^{M-1}$, that is $j_{k}=k$ for $k\in\llbracket 0,\eta\rrbracket$. We will show that we can sort one extra element.
@@
Data structuring: Let $k_{1},\,k_{2}\in\llbracket\eta,M-1\rrbracket$ be such that $z_{k_{1},j_{\eta+2}-1}^{\ell}$ and $z_{k_{2},j_{\eta+2}+1}^{\ell}$ are the left and right neighborhood points of $z_{\eta+2,j_{\eta+2}}^{\ell}$, respectively (if $z_{k_{2},j_{\eta+2}+1}^{\ell}$ does not exist, it is enough to take $z_{k_{2},j_{\eta+2}+1}^{\ell}:=z_{k_{1},j_{\eta+2}}^{\ell}+1$). Thus, consider the parameters
$w_{\ell+1}^{1}$ $=1,\quad w_{\ell+1}^{2}=-1,\quad b_{\ell+1}^{1}=-\left(\frac{z_{k_{2},j_{\eta+2}+1}^{\ell}+z_{\eta+2,j_{\eta+2}}^{\ell}}{2}\right),$
$\qquad\text{and}\quad b_{\ell+1}^{2}=\frac{z_{\eta+2,j_{\eta+2}}^{\ell}+z_{k_{1},j_{\eta+2}-1}^{\ell}}{2}.$
Define $W_{\ell+1}=(w_{\ell+1}^{1},w_{\ell+1}^{2})^{\top}$ and $b_{\ell+1}=(b_{\ell+1}^{1},b_{\ell+1}^{2})^{\top}$ and set
$z^{\ell+1}_{k}=\bm{\sigma}(W_{\ell+1}z^{\ell}_{k}+b_{\ell+1}).$
We obtain that $z^{\ell+1}_{\eta+2}=(0,0)$ and $z^{\ell+1}_{\eta+1}=(a,0)$, for some $a>0$, is the farthest point in the $x-$axis. Moreover, for a decreasing sequence of positive numbers $\{a_{k}\}_{k=0}^{M-1}$, we deduce
$z^{\ell+1}_{k}=(0,a_{k}),\quad\text{for all }k\in\llbracket 0,k_{1}\rrbracket.$ (4.18)
Projection: Let $w_{\ell+2}\in\mathbb{R}^{2}$ and $b_{\ell+2}\in\mathbb{R}$ be such that
$0\leq w_{\ell+2}\cdot z_{k}^{\ell+1}+b_{\ell+2}<w_{\ell+2}\cdot z_{k+1}^{\ell+1}+b_{\ell+2}\quad\text{for all }k\in\llbracket 0,\eta\rrbracket,$ (4.19)
$w_{\ell+2}\cdot z_{k}^{\ell+1}+b_{\ell+2}<w_{\ell+2}\cdot z_{\eta+2}^{\ell+1}+b_{\ell+2}\quad\text{for all }k\in\llbracket 0,M-1\rrbracket.$ (4.20)
The assumptions about $W_{\ell+2}$ mentioned above are not restrictive. Condition (4.19) is feasible, as shown by (4.18), where the sequence $z_{k}^{\ell+1}$ is arranged on the $y-$axis for $k\in\llbracket 0,M-2\rrbracket$ as a decreasing sequence. Furthermore, (4.20) is achievable by selecting the appropriate slope of the hyperplane defined by the parameters (see Remark 4.1). We then set $z_{k,j}^{\ell+2}=\sigma(w_{\ell+2}\cdot z_{k}^{\ell+1}+b_{\ell+2}),$ and, by construction, we have that $j_{k}=k$ for $k\in\llbracket 1,\eta+1\rrbracket$ and $j_{\eta+2}=M-1$, concluding the induction.
@@ 4.4. Mapping to the respective labels
$\phi^{L_{2}}(x_{i})=\phi(\mathcal{W}^{L_{2}},\mathcal{B}^{L_{2}},z_{k})=\xi_{k},\quad\text{for all }k\in\llbracket 0,M-1\rrbracket.$ (4.21)
@@
$\phi^{L_{3}}(x_{i})=\phi(\mathcal{W}^{L_{3}},\mathcal{B}^{L_{3}},\xi_{k})=k,\quad\text{for all }k\in\llbracket 0,M-1\rrbracket.$ (4.22)
@@
(2) Inductive Step: Consider $\eta\in\llbracket 0,M-1\rrbracket$ and define $l:=1+2(\eta-2)$. Let us assume that $\xi_{k}^{l}=k$ for all $k\in\llbracket 0,\eta\rrbracket$. We will show that there exist parameters such that $\xi_{k}^{l+2}=k$ for all $k\in\llbracket 0,\eta+1\rrbracket$. We will proceed in two steps: data structuring and data projection.
@@
$w_{l+1}^{1}=\frac{2}{\xi_{\eta+1}^{l}-\xi_{\eta}^{l}},\quad w_{l+1}^{2}=1,\quad b_{l+1}^{1}=-\left(\frac{\xi_{\eta+1}^{l}+\xi_{\eta}^{l}}{\xi_{\eta+1}^{l}-\xi_{\eta}^{l}}\right),\quad\text{and}\quad b_{l+1}^{2}=0.$
Define $W_{l+1}=(w_{l+1}^{1},w_{l+1}^{2})^{\top}$ and $b_{l+1}=(b_{l+1}^{1},b_{l+1}^{2})^{\top}$, and consider
$\xi_{k}^{l+1}=\bm{\sigma}(W_{l+1}\xi_{k}^{l}+b_{l+1})$
so that $\xi_{\eta+1}^{l+1}=(1,\xi_{\eta+1}^{l})$ and
$\xi_{k}^{l+1}=(0,k),\quad\text{for all }k\in\llbracket 0,\eta\rrbracket,$
$\xi_{k}^{l+1}=(a_{k},c_{k}),\quad\text{for all }k\in\llbracket\eta+2,M-1\rrbracket,$
where $\{a_{k}\}_{k=\eta+2}^{M-1}$ and $\{c_{k}\}_{k=\eta+2}^{M-1}$ are two sequences of strictly increasing numbers such that $a_{k}>1$ and $c_{k}>\xi_{\eta+1}^{l}$ for all $k\geq\eta+2$.
Data projection. Finally, define $\xi_{k}^{l+2}=\sigma(w_{l+2}\cdot\xi_{k}^{l+1}+b_{l+2})$ with
$w_{l+2}=((\eta+1)-\xi_{\eta+1}^{l},1),\quad\text{and }\quad b_{l+2}=0$
so that $\xi_{k}^{l+2}=k$ for all $k\in\llbracket 0,\eta+1\rrbracket$.
@@
Summarizing, the input-output map $\phi^{L}$ of Theorem 1.1, for $N$ points with $M$ classes, is given by the composition of the mappings $\phi_{i}^{L_{i}}$ given by (4.2), (4.13), (4.14) and (4.22) respectively, i.e.,
$\phi^{L}=(\phi^{L_{3}}\circ\phi^{L_{2}}\circ\phi^{L_{1}}\circ\phi^{L_{0}}),$
@@
Proof of Corollary 1.1: Now, we continue with the proof of Corollary 1.1. To this end, let us recall that we have assumed $\{x_{i},y_{i}\}\subset B_{R_{x}}^{d}(0)\times B_{R_{y}}^{1}(0)$, with $R_{x},R_{y}>0$. We proceed step by step to estimate the norms.
Preconditioning of the data: Denote by $(\mathcal{W}_{1},\mathcal{B}_{1})=(W_{1},b_{1})$ the parameters in this step. Recall that $\|w_{1}\|_{2}=1$ and we take $b_{1}=2R_{x}$, so that
$\|W_{1}\|_{F}=\|w_{1}\|_{2}=1,\quad\|b_{1}\|_{2}=2R_{x},\quad\|W_{1}\|_{\infty}\leq 1,\quad\|b_{1}\|_{\infty}=2R_{x}.$
Hence, we have that
$\left|\!\left|\!\left|(\mathcal{W}_{1},\mathcal{B}_{1})\right|\!\right|\!\right|_{2}^{2}$ $=\sum_{j=1}^{1}\bigl{(}\|W_{j}\|_{F}^{2}+\|b_{j}\|_{2}^{2}\bigr{)}=1^{2}+(2R_{x})^{2}=1+4R_{x}^{2},$ (4.23)
and for the $l^{\infty}-$ norm we get
$\left|\!\left|\!\left|(\mathcal{W}_{1},\mathcal{B}_{1})\right|\!\right|\!\right|_{\infty}$ $=\max_{j\in\{1\}}\{\|W_{j}\|_{\infty},\|b_{j}\|_{\infty}\}=\max\{1,2R_{x}\}=2R_{x}.$ (4.24)
Compression step: The total depth of the network is $L_{2}=2N$. For each $j\in\{0,\dots,N-1\}$, we first analyze the data structuring layer ($W_{2j}$, $b_{2j}$):
$W_{2j}=\begin{pmatrix}1\\ -1\end{pmatrix},\quad\text{so that}\quad\|W_{2j}\|_{F}^{2}=2,\quad\|b_{2j}\|_{2}^{2}\leq 2M_{1}^{2}.$
Now, for the Compression layer ($W_{2j+1}$, $b_{2j+1}$):
$W_{2j+1}=\begin{pmatrix}A_{j}&1\\ -B_{j}&-1\end{pmatrix},\quad\text{with}\quad A_{j},B_{j}\leq 1,\quad\|W_{2j+1}\|_{F}^{2}\leq 4,\quad\|b_{2j+1}\|_{2}^{2}\leq 2M_{1}^{2}.$
Here $M_{1}=\max_{i}\{\|W_{1}x_{1}+b_{1}\|_{\infty}\}$ and therefore $M_{1}\leq 3R_{x}$. To compute the full norm in this step, we sum over $j$ from 00 to $N-1$, obtaining
$\left|\!\left|\!\left|(\mathcal{W}_{2},\mathcal{B}_{2})\right|\!\right|\!\right|_{2}^{2}$ $=\sum_{j=0}^{N-1}\left(\|W_{2j}\|_{F}^{2}+\|b_{2j}\|_{2}^{2}+\|W_{2j+1}\|_{F}^{2}+\|b_{2j+1}\|_{2}^{2}\right)$
$=N(2+2M_{1}^{2})+N(4+2M_{1}^{2})=N(6+4M_{1}^{2}).$ (4.25)
For the $\ell^{\infty}$-norm we have that $\|W_{j}\|_{\infty}\leq 1,$ and $\|b_{j}\|_{\infty}\leq M_{1},$ thus
${\left|\!\left|\!\left|(\mathcal{W}_{2},\mathcal{B}_{2})\right|\!\right|\!\right|_{\infty}\leq 3R_{x}}.$ (4.26)
Data sorting: Let $R_{z}:=\max_{k}\|{z_{k}}\|_{\infty}$ being $\{z_{k}\}$ the output from the compression step. The network depth is $L_{3}=2M+1$. We consider projection layers indexed by $j\in\{0,2,\dots,2M\}$ and data structuring layers indexed by $j\in\{1,3,\dots,2M-1\}$. For each projection layer $W_{2j}$ and $b_{2j}$:
$\|W_{2j}\|_{F}^{2}=1,\quad\|b_{2j}\|_{2}^{2}\leq R_{z}^{2}.$
There are $M+1$ such layers. For each structuring layer $W_{2j+1}$ and $b_{2j+1}$:
$\|W_{2j+1}\|_{F}^{2}=2,\quad\|b_{2j+1}\|_{2}^{2}\leq 2R_{z}^{2}.$
There are $M$ such layers. The total norm is
$\left|\!\left|\!\left|(\mathcal{W}_{3},\mathcal{B}_{3})\right|\!\right|\!\right|_{2}^{2}$ $=\sum_{j=0}^{M}\left(\|W_{2j}\|_{F}^{2}+\|b_{2j}\|_{2}^{2}\right)+\sum_{j=0}^{M-1}\left(\|W_{2j+1}\|_{F}^{2}+\|b_{2j+1}\|_{2}^{2}\right)$
$=(M+1)(1+R_{z}^{2})+M(2+2R_{z}^{2})=(3M+1)(1+R_{z}^{2}).$
To estimate $R_{z}$, let $z_{i}^{(0)}=x_{i}^{1}$ be the output of Step 1, and assume $\|z_{i}^{(0)}\|_{\infty}\leq M_{1}$ for all $i$. Then each iteration in Step 2 satisfies:
$\|z_{i}^{(\ell)}\|_{\infty}\leq\|W_{\ell}\|_{\infty}\cdot\|z_{i}^{(\ell-1)}\|_{\infty}+\|b_{\ell}\|_{\infty}\leq\|z_{i}^{(\ell-1)}\|_{\infty}+M_{1}.$
By induction over $2N$ layers, we obtain:
$R_{z}:=\max_{i}\|z_{i}^{(2N)}\|_{\infty}\leq(2N+1)M_{1}\leq 2R_{x}(2N+1).$
Using the bound for $R_{z}$ derived above,
${\left|\!\left|\!\left|(\mathcal{W}_{3},\mathcal{B}_{3})\right|\!\right|\!\right|_{2}^{2}\leq{(3M+1)(1+4R_{x}^{2}(2N+1)^{2})}}.$ (4.27)
For the $\ell^{\infty}$-norm, since all parameters are bounded by $R_{z}$, we deduce
${\left|\!\left|\!\left|(\mathcal{W}_{3},\mathcal{B}_{3})\right|\!\right|\!\right|_{\infty}\leq 2R_{x}(2N+1)}.$ (4.28)
Mapping to the labels: We denote by $(\mathcal{W}_{4},\mathcal{B}_{4})$ the collection of parameters used in the final step of the construction. Recall that the total number of layers in this step is $L_{4}=2M-3$. The first three layers implement the initial projection and positioning of the first three values. All parameters involved in these layers are explicitly defined and uniformly bounded in terms of the geometry of the values $\{\xi_{k}\}_{k=0}^{M-1}$. Therefore, there exists a constant $C_{\xi}>0$, depending only on the relative distance of the data $\{x_{i}\}_{i=1}^{N}$, such that
$\|W_{j}\|_{F}^{2}+\|b_{j}\|_{2}^{2}\leq C_{\xi}(1+R_{y}^{2}),\quad\|W_{j}\|_{\infty},\|b_{j}\|_{\infty}\leq\max\{1,C_{\xi}\},\quad\text{for}j=1,2,3.$
In the inductive step, we have that for each $\eta\in\llbracket 2,M-2\rrbracket$, two layers are required to map $\xi_{\eta+1}\mapsto\eta+1$. This is done with the Data structuring layer, for which we have Parameters satisfy again
$\|W_{j}\|_{F}^{2}+\|b_{j}\|_{2}^{2}\leq C_{\xi}\,(1+R_{y}^{2}),\quad\|W_{j}\|_{\infty},\|b_{j}\|_{\infty}\leq\max\{1,\,C_{\xi}\}.$
Then we apply the Projection layer with weight $w=(\eta+1-\xi_{\eta+1},1)$, and bias 00, and hence
$\|W_{j}\|_{F}^{2}=(\eta+1-\xi_{\eta+1})^{2}+1\leq(M+R_{y})^{2},\quad\|b_{j}\|_{2}^{2}=0,\quad\|W_{j}\|_{\infty}$ $\leq M+R_{y},\quad\|b_{j}\|_{\infty}=0.$
Then, putting all together, we have three initial layers plus $(M-3)$ inductive steps. Therefore, we deduce
$\left|\!\left|\!\left|(\mathcal{W}_{4},\mathcal{B}_{4})\right|\!\right|\!\right|_{2}^{2}$ $=3\,C_{\xi}\,(1+R_{y}^{2})+(M-3)\bigl{[}C_{\xi}\,(1+R_{y}^{2})+(M+R_{y})^{2}\bigr{]}$
$=M\,C_{\xi}\,(1+R_{y}^{2})+(M-3)\,(M+R_{y})^{2}.$ (4.29)
Analogously for the $\ell^{\infty}$ norm, taking the maximum over all layers yields
${\left|\!\left|\!\left|(\mathcal{W}_{4},\mathcal{B}_{4})\right|\!\right|\!\right|_{\infty}\leq\max\bigl{\{}\max\{1,C_{\xi}\},\,M+R_{y}\bigr{\}}.}$ (4.30)
Finally, combining estimation (4.23)-(4.4)-(4.27)-(4.4), we deduce that
[…]
$\quad+\bigl{[}M\,C_{\xi}(1+R_{y}^{2})+(M-3)(M+R_{y})^{2}\bigr{]},$
and for the $l^{\infty}-$norm we combine estimations (4.24)-(4.26)-(4.28)-(4.30) to obtain
[…]
In particular, we can guarantee the existence of a constant $C>0$ independent of $N,\,M,\,R_{x},\,R_{y}$ such that
$\left|\!\left|\!\left|(\mathcal{W},\mathcal{B})\right|\!\right|\!\right|_{2}$ $\leq C(1+R_{x}\sqrt{N}+R_{x}N\sqrt{M}+R_{y}M),$
and
$\left|\!\left|\!\left|(\mathcal{W},\mathcal{B})\right|\!\right|\!\right|_{\infty}$ $\leq C\bigl{(}R_{x}N+M+R_{y}\bigr{)}.$